OpenStudy (anonymous):

Solve the equation. 4x + 3.6 + x = 1.2 A. x = -0.6 B. x = 1.2 C. x = -0.48 D. x = 0.96

OpenStudy (nenadmatematika):

c

OpenStudy (anonymous):

4x + 3.6 + x = 1.2 5x = 1.2 - 3.6 5x = -2.4 x = -2.4 / 5 x = =0.48. =)

OpenStudy (anonymous):

-.48